# Robust Deterministic Policy Gradient in Zero-Sum Games for Disturbance Attenuation

### v5

@@ Abstract
Practical control systems pose significant challenges in identifying optimal control policies due to uncertainties in the system model and external disturbances. While $H_{\infty}$ control techniques are commonly used to design robust controllers that mitigate the effects of disturbances, these methods often require complex and computationally intensive calculations. To address this issue, this paper proposes a reinforcement learning algorithm called robust deterministic policy gradient (RDPG), which formulates the $H_{\infty}$ control problem as a two-player zero-sum dynamic game. In this formulation, one player (the user) aims to minimize the cost, while the other player (the adversary) seeks to maximize it. We then employ deterministic policy gradient (DPG) and its deep reinforcement learning counterpart to train a robust control policy with effective disturbance attenuation. In particular, for practical implementation, we introduce an algorithm called robust deep deterministic policy gradient (RDDPG), which employs a deep neural network architecture and integrates techniques from the twin-delayed deep deterministic policy gradient (TD3) to enhance stability and learning efficiency. To evaluate the proposed algorithm, we implement it on an unmanned aerial vehicle (UAV) tasked with following a predefined path in a disturbance-prone environment. The experimental results demonstrate that the proposed method outperforms other control approaches in terms of robustness against disturbances, enabling precise real-time tracking of moving targets even under severe disturbance conditions.
@@
Deep reinforcement learning, Two players zero-sum game, Unmanned aerial vehicles
@@ 1 Introduction
In recent years, deep reinforcement learning (DRL) has shown remarkable success in finding optimal policies across various domains, including games [1, 2] and control tasks [3, 4]. In particular, various model-free DRL algorithms, such as, deep deterministic policy gradient (DDPG) [3], twin delayed DDPG (TD3) [5], soft actor-critic (SAC) [6], and proximal policy optimization (PPO) [7], have demonstrated the ability to generate effective control inputs without requiring explicit system models. However, a critical challenge arises from the discrepancy between the simulation environments used for training and the real-world environment [8, 9, 10]. Therefore, developing DRL policies that are robust to unseen disturbances is crucial for ensuring reliability in real-world applications. To address this challenge, in this paper, we introduce adversarial learning, where the agent is trained in the presence of disturbances. By exposing the policy to worst-case scenarios during training, the agent can improve its robustness and generalization to unseen environments.
@@
In this paper, we present the so-called robust deterministic policy gradeint (RDPG) algorithm which enhances the robustness of deterministic policy gradient (DPG) algorithm [11] from the concept of the $H_{\infty}$ control and the two-player zero-sum dynamic game perspectives. The $H_{\infty}$ control [12, 13, 14] method is a robust control technique that aims to design an optimal controller that minimizes the impact of disturbances on system performance. A well-established result in control theory shows that, under certain assumptions, the $H_{\infty}$ control problem can be reformulated as a two-player zero-sum dynamic game, or equivalently, a min-max optimization problem [15]. In this game-theoretic formulation, the controller (first player) seeks to minimize a given cost function, while the adversary (second player) simultaneously tries to maximize it by injecting disturbances [15].
@@
Building upon this theoretical foundation, we transform the $H_{\infty}$ control problem into a two-player zero-sum dynamic game between the user (the controller) and the adversarial agent. We then propose a reinforcement learning approach based on the DPG method to simultaneously train both agents. Each agent learns its own policy: the user aims to minimize the objective function, while the adversary aims to maximize it. As a result of this zero-sum game framework, the user learns a control policy that is inherently robust against worst-case disturbances. Furthermore, we extend this idea to high-dimensional continuous control settings by incorporating it into a deep reinforcement learning framework. The resulting algorithm, named robust deep deterministic policy gradient (RDDPG), combines the core idea of RDPG with TD3 [5].
@@
Additionally, we aim to develop a robust control strategy for quadrotors, a widely used unmanned aerial vehicle (UAV) in various applications. The ability to maintain robust control of quadrotors is particularly important in dynamic and uncertain environments, where tasks such as tracking moving targets or following predefined waypoints are required. We demonstrate that the proposed RDDPG method effectively determines optimal control inputs in the presence of external disturbances. Furthermore, it outperforms existing DRL-based control approaches by achieving lower cost under severe disturbance conditions.
@@
    We apply the proposed RDDPG algorithm to the robust tracking control of quadrotors and conduct comprehensive numerical simulations. The results demonstrate that RDDPG significantly outperforms several state-of-the-art DRL methods, including DDPG, SAC, and TD3.
@@ 2 Related works
The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems [14, 16, 17, 18]. However, these methods require precise model linearization and incur high computational costs due to the need to solve Hamilton-Jacobi-Isaacs (HJI) equations. To address these limitations, alternative approaches leverage deep reinforcement learning (DRL) to determine optimal control policies without the need for solving nonlinear HJI equations.
Recent efforts have extended the ideas of $H_{\infty}$ control and two-player zero-sum games into the DRL framework to enhance robustness under various scenarios. For instance, Deshpande et al. [19] developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks. Their method primarily addresses model uncertainties rather than external disturbances by adopting an action-robust reinforcement learning framework [9], where an adversary injects noise directly into the agent’s actions to minimize the agent’s reward. In contrast, our method explicitly introduces external disturbances into the environment dynamics and formulates the learning process as a two-player zero-sum game between the agent and the adversary. By doing so, our approach enables the agent to learn a control policy that is robust to worst-case disturbance scenarios, rather than merely compensating for action-level noise.  Long et al. [10] proposed a novel approach that models robust locomotion learning as an interaction between the locomotion policy and a learnable disturbance generator. The disturbance is conditioned on the robot’s state and produces appropriate external forces to facilitate policy learning. To address modeling errors and discrepancies between training and testing conditions, robust adversarial reinforcement learning (RARL) [8] was proposed. RARL trains the agent to perform effectively in the presence of a destabilizing adversary that applies disturbance forces to the system. The agent and the adversary are concurrently trained using the trust region policy optimization (TRPO) algorithm, with the agent learning to maximize the expected reward while the adversary learns to minimize it. Although the underlying concept is similar, their approach differs significantly from ours. Both methods adopt stochastic policy frameworks, such as PPO and TRPO, and they do not impose explicit constraints on the adversary, which may lead to the generation of excessively strong or unrealistic disturbances during training. In contrast, our method is based on deterministic policy gradients and explicitly incorporates the $H_{\infty}$ control perspective into a two-player zero-sum game formulation, enabling more stable and practical learning in continuous control tasks.
@@ 3.1 $H_{\infty}$ Control
${J^{\pi}}:=\mathop{\sup}\limits_{{\bf w}_{0:\infty}\neq 0}{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi}\right]$
@@
3.2 Two-player zero-sum dynamic game perspective
According to [15], the suboptimal $H_{\infty}$ control problem can be equivalently viewed as solving a zero-sum dynamic game under certain assumptions. In particular, we consider two decision making agents called the user and the adversary, respectively. They sequentially take control actions $u_{k}$ (user) and $w_{k}$ (adversary) to minimize and maximize cumulative discounted costs, respectively, of the form
${\mathbb{E}}\left[{\sum_{k=0}^{\infty}{\gamma^{k}c(x_{k},u_{k},w_{k})}}\right]$
where $\gamma\in(0,1]$ is the discount factor, $u_{k}$ is the control input of the user, and $w_{k}$ is the control input of the adversary. The two agents have opposing objectives. The user’s primary objective entails minimizing the costs, while the adversary strives to hinder the user’s progress by maximizing the costs. There exist two categories of the two-player dynamic games, the alternating two-player dynamic game and simultaneous two-player dynamic game. In the alternating two-player dynamic game, two agents engaged in decision making take turns in selecting actions to maximize and minimize the cost, respectively. On the other hand, in the simultaneous two-player dynamic game, the two agents take actions simultaneously to maximize and minimize the cost. In this paper, we mainly focus on the simultaneous two-player dynamic game.
@@ 3.2 Two-player zero-sum dynamic game perspective
The corresponding cost function is defined as
$J^{\pi,\mu}:={\mathbb{E}}\left[\left.\sum\limits_{k=0}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k})\right|\pi,\mu\right]$ (4)
The goal of the dynamic game is to find a saddle-point pair of equilibrium policies $(\pi,\mu)$ (if exists) for which
@@
By the min-max theorem, this also implies that
${\max_{\mu\in{\rm M}}}{\min_{\pi\in\Pi}}{J^{\pi,\mu}}={\min_{\pi\in\Pi}}{\max_{\mu\in{\rm M}}}{J^{\pi,\mu}}$
The optimal cost is now defined as
${J^{*}}:={J^{{\pi^{*}},{\mu^{*}}}}$
To show a connection between the dynamic game framework and the $H_{\infty}$ control problem, let us consider the specific per-step cost function
$c(x,u,w)=\left\|{g(x,u,w)}\right\|_{2}^{2}-{\eta^{2}}\left\|w\right\|_{2}^{2}.$
@@
${J^{\pi,\mu}}={\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi,\mu}\right]$
@@
which is structurally very similar to the policy with the $H_{\infty}$ performance in the previous subsection. It is known that under some special conditions such as the linearity, the $H_{\infty}$ suboptimal control policy and the saddle-point policy are equivalent when $J^{*}\leq 0$.
@@ 3.3 Bellman equations
${V^{\pi,\mu}}(x):={\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k})}\right|\pi,\mu,x_{t}=x}\right]$
@@
$V^{\pi,\mu}(x)=c(x,\pi(x),\mu(x))+\gamma{\mathbb{E}}[{V^{\pi,\mu}}(x^{\prime})|\pi,\mu]$
@@
${Q^{\pi,\mu}}(x,u,w):={\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}{\gamma^{k}c({x_{k}},{u_{k}},{w_{k}})}}\right|{x_{t}}=x,{u_{t}}=u,{w_{t}}=w,\pi,\mu}\right]$
@@
${Q^{\pi,\mu}}(x,u,w)=c(x,u,w)+\gamma{\mathbb{E}}[{Q^{\pi,\mu}}(x^{\prime},u^{\prime},w^{\prime})|\pi,\mu]$
@@
${Q^{*}}(x,u,w)=c(x,u,w)+\gamma{\mathbb{E}}[{Q^{*}}(x^{\prime},u^{\prime},w^{\prime})|{\pi^{*}},{\mu^{*}}]$
@@ 4 Proposed method
In this section, we will present the proposed method, robust deterministic policy gradient (RDPG) and its deep reinforcement learning version, robust deep deterministic policy gradient (RDDPG).
Figure 1 illustrates the architecture of the proposed method, RDDPG, based on a two-player zero-sum dynamic game. In this method, two policy, user policy $\pi$ and adversary policy $\mu$, interact within the environment. The user policy generates the action $u_{t}$, while the adversary policy generates the disturbance $w_{t}$, both based on the current state $x_{t}$. These inputs are fed into the environment, which returns the next state $x_{t+1}$ and the cost $c_{t}$ based on the dynamics of the environment. The critic is updated based on $\{x_{t},u_{t},w_{t},x_{t+1},c_{t}\}$ and evaluates the joint performance of the user and adversary. The policy of user and adversary are updated by deterministic policy gradient method to minimize and maximize the objective function $J^{\pi,\mu}$. The detailed procedure is described in the following sections.
Figure 1: Overview of RDDPG. Based on the two-player zero-sum dynamic game, the user policy $\pi$ seeks to minimize the objective function $J^{\pi,\mu}$, while the adversary policy $\mu$ attempts to maximize it by injecting disturbances into the environment.
@@ 4.1 Robust deterministic policy gradient
To apply deterministic policy gradient (DPG) [11] to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
@@
where $\alpha_{\theta}>0$ and $\alpha_{\phi}>0$ are step-sizes. According to [11], the deterministic gradients can be obtained using the following theorem.
Theorem 1.
@@
The proof is a simple extension of the DPG theorem in [11], so it is omitted here. A reinforcement learning counterpart to DPG can be obtained by using the samples of the gradients. Next, we will consider the following per-step cost function:
$c(x,u,w)=\tilde{c}(x,u,w)-\eta^{2}\left\|w\right\|_{2}^{2}$
where $\tilde{c}:{\mathbb{R}}^{p}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\times{\mathbb{R}}^{p}\to{\mathbb{R}}$ is the cost function for the user. Note that the cost function $\tilde{c}$ can be set arbitrarily and is more general than the quadratic cost in the $H_{\infty}$ control problem. This is because the $H_{\infty}$ control problem can be recovered by setting
$\tilde{c}(x,u,w)=\left\|{g(x,u,w,v)}\right\|_{2}^{2}$
In this way, we can consider the two-player zero-sum dynamic game as an extension of the usual reinforcement learning tasks by considering robustness against the adversarial disturbance. The additional disturbance quadratic cost ${\eta^{2}}w^{T}w$ in the cost function discourages the adversary from generating excessively large disturbances, effectively regularizing its behavior and ensuring that the learned policies are robust against realistic worst-case scenarios rather than arbitrary or unbounded inputs.
@@
$\frac{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{\gamma^{k}\tilde{c}({x_{k}},{u_{k}},{w_{k}})}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}\gamma^{k}\left\|w_{k}\right\|_{2}^{2}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}\leq{\eta^{2}}$

### v6

# Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control

@@ Abstract
This paper presents a robust reinforcement learning algorithm called robust deterministic policy gradient (RDPG), which reformulates the $H_{\infty}$ control problem as a two-player zero-sum dynamic game between a user and an adversary. The method combines deterministic policy gradients with deep reinforcement learning to train a robust policy that attenuates disturbances efficiently. A practical variant, robust deep deterministic policy gradient (RDDPG), integrates twin-delayed updates for stability and sample efficiency. Experiments on an unmanned aerial vehicle demonstrate superior robustness and tracking accuracy under severe disturbance conditions.
@@
Reinforcement learning and deep learning in control, Robust control applications, Trajectory tracking and path following for AVs, AI and learning-based control for automotive systems.
thanks: This work was supported by the Institute of Information Communications Technology Planning Evaluation (IITP) funded by the Korea government under Grant 2022-0-00469.
@@ 1 Introduction
In recent years, deep reinforcement learning (DRL) has shown remarkable success in finding optimal control policies across various domains, including games (RL1; DQN) and control tasks (DDPG; RL2), where various model-free algorithms have demonstrated the ability to generate effective control inputs without requiring explicit system models (DDPG; TD3; SAC; PPO). However, in real-world applications, physical systems are inevitably subject to external disturbances, unmodeled dynamics, and variations in system parameters. These factors often lead to significant discrepancies between the training simulation environments and the conditions encountered in practice (RARL; AR-MDP; HRL). Consequently, policies trained in simulation frequently suffer from performance degradation or instability when deployed on real systems. This highlights that developing DRL policies that remain robust under unseen disturbances and model uncertainties is essential for ensuring reliability in practical applications.
@@
In classical control theory, the $H_{\infty}$ control framework (H1; H2; Hinf1) provides a principled approach for designing controllers that minimize the worst-case impact of disturbances on system performance. A well-known result shows that, under certain assumptions, the $H_{\infty}$ control problem can be reformulated as a two-player zero-sum dynamic game (TZDG), or equivalently, a min–max optimization problem (Hbook). In this game-theoretic formulation, the controller (first player) seeks to minimize a given cost function, while the adversary (second player) simultaneously tries to maximize it by injecting disturbances (Hbook).
@@
Building upon this theoretical foundation, we introduce the robust deterministic policy gradient (RDPG) algorithm, which enhances the robustness of the deterministic policy gradient (DPG) method (DPG) by transforming the $H_{\infty}$ control formulation into a TZDG between the user (controller) and an adversarial agent. We develop a reinforcement learning framework in which both agents are trained simultaneously: the user aims to minimize the objective, while the adversary attempts to maximize it. This min–max interplay naturally encourages the user to acquire a policy that is robust to worst-case disturbances. Furthermore, we extend this idea to high-dimensional continuous control settings by incorporating it into a deep reinforcement learning framework. The resulting algorithm, named robust deep deterministic policy gradient (RDDPG), combines the core idea of RDPG with TD3 (TD3).
@@
Additionally, we aim to develop a robust control strategy for quadrotors, a widely used unmanned aerial vehicle (UAV) in various applications. The ability to maintain robust control of quadrotors is particularly important in dynamic and uncertain environments, where tasks such as tracking moving targets or following predefined waypoints are required. To this end, we demonstrate that the proposed RDDPG method is capable of generating optimal control inputs even in the presence of external disturbances. Experimental results show that RDDPG achieves lower costs and exhibits stronger robustness under disturbance conditions compared to existing DRL-based control approaches.
@@
    We apply the proposed RDDPG algorithm to the tracking control of quadrotor and conduct comprehensive numerical simulations. The results demonstrate that RDDPG significantly outperforms several state-of-the-art DRL and robust RL methods.
@@ 2 Related works
The $H_{\infty}$ control methods are effective in designing robust controllers against disturbances and have demonstrated satisfactory performance in nonlinear dynamic systems (Hinf1; Hinf2; Hinf3; Hinf5). However, these methods require precise model linearization and incur high computational costs due to the need to solve Hamilton-Jacobi-Isaacs (HJI) equations. To address these limitations, alternative approaches leverage deep reinforcement learning (DRL) to determine optimal control policies without the need for solving nonlinear HJI equations.
Recent efforts have extended the ideas of $H_{\infty}$ control and two-player zero-sum games into the DRL framework to enhance robustness under various scenarios. For instance, RRLforquad developed a robust DRL approach to bridge the gap between different environments for quadcopter control tasks. Their method primarily addresses model uncertainties rather than external disturbances by adopting an action-robust reinforcement learning framework (AR-MDP), where an adversary injects noise directly into the agent’s actions to minimize the agent’s reward. In contrast, our method explicitly introduces external disturbances into the environment dynamics and formulates the learning process as a two-player zero-sum game between the agent and the adversary. By doing so, our approach enables the agent to learn a control policy that is robust to worst-case disturbance scenarios, rather than merely compensating for action-level noise.  HRL proposed a novel approach that models robust locomotion learning as an interaction between the locomotion policy and a learnable disturbance generator. The disturbance is conditioned on the robot’s state and produces appropriate external forces to facilitate policy learning. To address modeling errors and discrepancies between training and testing conditions, robust adversarial reinforcement learning (RARL) (RARL) was proposed. RARL trains the agent to perform effectively in the presence of a destabilizing adversary that applies disturbance forces to the system. The agent and the adversary are concurrently trained using the trust region policy optimization (TRPO) algorithm, with the agent learning to maximize the expected reward while the adversary learns to minimize it. Although the underlying concept is similar, their approach differs significantly from ours. Both methods adopt stochastic policy frameworks and they do not impose explicit constraints on the adversary, which may lead to the generation of excessively strong or unrealistic disturbances during training. In contrast, our method is based on deterministic policy gradients and explicitly incorporates the $H_{\infty}$ control perspective into a two-player zero-sum game formulation, enabling more stable and practical learning in continuous control tasks.
@@ 3.1 $H_{\infty}$ Control
${J^{\pi}}:=\mathop{\sup}\limits_{{\bf w}_{0:\infty}\neq 0}{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|y_{k}\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi}\right]$ (3)
@@
3.2 Two-player zero-sum dynamic game
According to (Hbook), the suboptimal $H_{\infty}$ control problem can be equivalently viewed as solving a two-player zero-sum dynamic game (TZDG) under certain assumptions. In particular, we consider two decision making agents called a user and an adversary, respectively. TZDG can be expressed as a tuple $(\mathcal{S},\mathcal{A}_{1},\mathcal{A}_{2},P,c,\gamma,x_{0})$. Here $\mathcal{S}=\mathbb{R}^{p}$ is the state space, $\mathcal{A}_{1}=\mathbb{R}^{m}$ and $\mathcal{A}_{2}=\mathbb{R}^{n}$ are the continuous action spaces for the user and adversary, respectively. The user selects an control input $u\in\mathcal{A}_{1}$ and the second agent selects disturbance $w\in\mathcal{A}_{2}$ at the current state $x\in\mathcal{S}$ simultaneously. Then the state transits to the next state $x^{\prime}\in\mathcal{S}$ with the state transition probability $P(x^{\prime}|x,u,w)$ and the cost $c\in\mathbb{R}$ is incurred by the cost function $c(x,u,w,x^{\prime}):\mathcal{S}\times\mathcal{A}_{1}\times\mathcal{A}_{2}\times\mathcal{S}\rightarrow\mathbb{R}$. The $\gamma\in(0,1]$ is the discounted factor and $x_{0}\in\mathcal{S}$ represents the initial state.
@@ 3.2 Two-player zero-sum dynamic game
The objective of the user and adversary are to minimize and maximize the cumulative discounted costs over infinite time horizon $J^{\pi,\mu}$ defined as
$J^{\pi,\mu}:={\mathbb{E}}\left[\left.\sum\limits_{k=0}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k},x_{k+1})\right|\pi,\mu\right]$
where $\mathbb{E}[\cdot|\pi,\mu]$ is an expectation conditioned on the two policies $\pi$ and $\mu$. The goal of the dynamic game is to find a saddle-point pair of equilibrium policies $(\pi,\mu)$ (if exists) for which
@@
perolat2015approximate; AR-MDP demonstrated that for a game with optimal equilibrium return $J^{*}$, there always exists the Nash equilibrium, and it is equivalent to the minimax solution,
${J^{*}}:={J^{{\pi^{*}},{\mu^{*}}}}={\max_{\mu\in{\rm M}}}{\min_{\pi\in\Pi}}{J^{\pi,\mu}}={\min_{\pi\in\Pi}}{\max_{\mu\in{\rm M}}}{J^{\pi,\mu}}$ (4)
To show a connection between this TZDG framework and the $H_{\infty}$ control problem, let us consider the specific per-step cost function
$c(x_{k},u_{k},w_{k},x_{k+1})=\left\|{g(x_{k},u_{k},w_{k},x_{k+1})}\right\|_{2}^{2}-{\eta^{2}}\left\|w_{k}\right\|_{2}^{2}.$
@@
${J^{\pi,\mu}}={\mathbb{E}}\left[\left.\sum\limits_{k=0}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k},x_{k+1})\right|\pi,\mu\right]$
$={\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{(\|g(x_{k},u_{k},w_{k},x_{k+1})\|_{2}^{2}-\eta^{2}\|w_{k}\|_{2}^{2})}}\right|\pi,\mu}\right].$ (5)
@@
which is structurally very similar to Equation (3), the policy with the $H_{\infty}$ performance in the previous subsection. It is known that under some special conditions such as the linearity, the $H_{\infty}$ suboptimal control policy and the saddle-point policy are equivalent when $J^{*}\leq 0$.
@@ 3.3 Bellman equations
${V^{\pi,\mu}}(x):={\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}\gamma^{k}c(x_{k},u_{k},w_{k},x_{k+1})}\right|\pi,\mu,x_{k}=x}\right]$
@@
$V^{\pi,\mu}(x)=c(x,\pi(x),\mu(x),x^{\prime})+\gamma{\mathbb{E}}[{V^{\pi,\mu}}(x^{\prime})|\pi,\mu]$
@@
${Q^{\pi,\mu}}(x,u,w):=$
${\mathbb{E}}\left[{\left.{\sum\limits_{k=t}^{\infty}{\gamma^{k}c(x_{k},u_{k},w_{k},x_{k+1})}}\right|{x_{k}}=x,{u_{k}}=u,{w_{k}}=w,\pi,\mu}\right]$
@@
${Q^{\pi,\mu}}(x,u,w)=c(x,u,w,x^{\prime})+\gamma{\mathbb{E}}[{Q^{\pi,\mu}}(x^{\prime},u^{\prime},w^{\prime})|\pi,\mu]$
@@
${Q^{*}}(x,u,w)=c(x,u,w,x^{\prime})+\gamma{\mathbb{E}}[{Q^{*}}(x^{\prime},u^{\prime},w^{\prime})|{\pi^{*}},{\mu^{*}}]$
@@ 4 Proposed method
In this section, we will present the proposed method, robust deterministic policy gradient (RDPG) and its deep reinforcement learning version, robust deep deterministic policy gradient (RDDPG). Fig. 1 illustrates the architecture of the proposed method, RDDPG, based on the TZDG. In this method, two policy, user policy $\pi_{\theta}$ and adversary policy $\mu_{\phi}$, interact within the environment. The user policy generates the action $u_{k}$, while the adversary policy generates the disturbance $w_{k}$, both based on the current state $x_{k}$. These inputs are fed into the environment, which returns the next state $x_{k+1}$ and the cost $c_{k+1}$ based on the dynamics of the environment. The critic is updated based on $x_{k},u_{k},w_{k},x_{k+1},c_{k+1}$} and evaluates the joint performance of the user and adversary. The policy of user and adversary are updated by deterministic policy gradient method to minimize and maximize the objective function $J^{\pi_{\theta},\mu_{\phi}}$. The detailed procedure is described in the following sections.
Figure 1: Overview of RDDPG. Based on the two-player zero-sum dynamic game, the user policy $\pi_{\theta}$ seeks to minimize the objective function $J^{\pi_{\theta},\mu_{\phi}}$, while the adversary policy $\mu_{\phi}$ attempts to maximize it by injecting disturbances into the environment.
@@ 4.1 Robust deterministic policy gradient
To apply deterministic policy gradient (DPG) (DPG) to the two-player zero-sum dynamic games, we consider the parameterized deterministic control policies
@@
where $\alpha_{\theta}>0$ and $\alpha_{\phi}>0$ are step-sizes. According to DPG, the deterministic gradients can be obtained using the following theorem.
Theorem 1
@@
The proof is a simple extension of the DPG theorem (DPG), so it is omitted here. A reinforcement learning counterpart to DPG can be obtained by using the samples of the gradients. Next, we will consider the following per-step cost function:
$c(x_{k},u_{k},w_{k},x_{k+1})=\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})-\eta^{2}\left\|w_{k}\right\|_{2}^{2}$
where $\tilde{c}:{\mathbb{R}}^{p}\times{\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\times{\mathbb{R}}^{p}\to{\mathbb{R}}$ is the cost function for the user. Note that the cost function $\tilde{c}$ can be set arbitrarily and is more general than the quadratic cost in the $H_{\infty}$ control problem. This is because the $H_{\infty}$ control problem can be recovered as shown Equation (3) and (5) by setting
$\tilde{c}(x_{k},u_{k},w_{k},x_{k+1})=\left\|{g(x_{k},u_{k},w_{k},x_{k+1})}\right\|_{2}^{2}$
In this way, we can consider the TZDG as an extension of the usual reinforcement learning tasks by considering robustness against the adversarial disturbance. The additional disturbance quadratic cost ${\eta^{2}}w^{T}w$ in the cost function discourages the adversary from generating excessively large disturbances, effectively regularizing its behavior and ensuring that the learned policies are robust against realistic worst-case scenarios rather than arbitrary or unbounded inputs.
@@
$\frac{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}{\gamma^{k}\tilde{c}({x_{k}},{u_{k}},{w_{k}},{x_{k+1}})}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}{{{\mathbb{E}}\left[{\left.{\sum\limits_{k=0}^{\infty}\gamma^{k}\left\|w_{k}\right\|_{2}^{2}}\right|{\pi_{\theta_{\varepsilon}^{*}}},{\mu_{\phi_{\varepsilon}^{*}}}}\right]}}\leq{\eta^{2}}$